A jet lands on an aircraft carrier at 140 mi/h ( ).
(a.) What is its acceleration if it stops in 2.0 s?
(b.) What is the displacement of the plane while it is stopping?
1
Expert's answer
2020-12-06T17:20:02-0500

a)


"a=\frac{v}{t}""a=\frac{63}{2}=31.5\frac{m}{s^2}"

b)



"s=0.5at^2""s=0.5(31.5)(2)^2=63\ m"